hartnn (hartnn):

2x^2-19x+24 @Laiken98 we are here, 2x^2-3x-16x+24 clear till here, right ?

hartnn (hartnn):

now factoring out x from 1st 2 terms,

OpenStudy (anonymous):

can i jump to the answer?

hartnn (hartnn):

2x^2 = 2*x*x -19x = -19*x the common term is 'x' what remains ?

hartnn (hartnn):

jump to answer ? why so hurry ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

quadratic or complete the square, or factor fractions your choice

OpenStudy (anonymous):

Well since the x's or gone only the numbers remain, correct?

hartnn (hartnn):

so, 2x^2-3x will be x (2x -3) see whether this makes sense ?

hartnn (hartnn):

yes, i made a typo, it should be 2x^2 = 2*x*x -3x = -3*x the common term is 'x'

OpenStudy (anonymous):

Oh, I understand now. Thank you!

hartnn (hartnn):

so, 2x^2 -3x -16x +24 = x (2x-3) -8 (2x-3) did u really understand this part ?

OpenStudy (anonymous):

I like your method of avoiding the quadratic hartnn